I need some help with this problem, it says to factor the trinomial, the trinomial is: 12x^2+10x-50, thanks.
Answer by checkley71(8403)

12X^2+10x-50=0
6x^2+5x-25=0
(3x-5)(2x+5)=0
3x-5=0
3x=5
x=5/3
2x+5=0
2x=-5
x=-5/2